A cable exerts a constant upward tension of magnitude 2.26  104 N on a 2.00  103 kg elevator as it rises through a vertical distance of 3.50 m.


(a) Find the work done by the tension force on the elevator (in J).

 

___ J


(b) Find the work done by the force of gravity on the elevator (in J).

 

 ___J


1
Expert's answer
2021-09-24T19:04:32-0400

Tension in the string "T = 2.26 \times 10^4\; N"

mass of the elevator "m = 2.00 \times 10^3 \;kg"

the rise in vertical distance S = 3.5 m

Tension force will act upwards on the elevator

displacement is also upwards

Work done "W = FScos \u03b8" here θ = 0

"W_1 = (2.26 \times 10^4 \; N) \times (3.5) \times cos 0 \\

W_1 = 7.91 \times 10^4 \;Joules"

Work done by the tension force on the elevator "W_1 = 7.91 \times 10^4 \;J"

Now,

gravitational force "F_g = mg"

"F_g = 2.00 \times 10^3\; kg \times 9.81 \\

F_g = 1.962 \times 10^4 \; kg"

Work done is "W = FS cos \u03b8" here θ is 180 (force and displacement are in the opposite direction)

"W = 1.962 \times 10^4 \; kg \times 3.5 \times (-1) \\

W = -6.87 \times 10^4 \; Joules"

Work done by the gravitational force on the elevator "W_1 = -6.87 \times 10^4 \;J"